2) When the real wage is above the level that equilibrates supply and demand:

Select one:

a. It creates a deadweight loss in the labor market.

b. the quantity of labor demanded exceeds the quantity supplied.

c. GDP definitely rises.

d. Interest rate rises.

3) If Central Bank A cares only about keeping the price level stable and Central Bank B cares only about keeping output at its natural level, then in response to an exogenous increase in the price of oil:

Select one:

a. both Central Bank A and Central Bank B should increase the quantity of money.

b. Central Bank A should increase the quantity of money, whereas Central Bank B should keep it stable.

c. Central Bank A should decrease the quantity of money, whereas Central Bank B should increase it.

d. both Central Bank A and Central Bank B should keep the quantity of money stable.

4) In a Keynesian Cross economy without the foreign sector, the marginal propensity to save is 0.2. Investment is 100; government expenditure is also 100. Taxes are 100. How much does total savings change if marginal propensity to save goes up from 0.2 to 0.3?

a. goes up by 90

b. goes up by 100

c. does not change

d. None of the above or cannot be determined without more information

5) Using the simple Keynesian Cross analysis, assume that the consumption function is given by C = 100 + 0.6(Y-T). If planned investment is 200 and T is 300, the level of G needed to make equilibrium Y equal 1,000 is (assume that net exports are zero):

Select one:

a. 60

b. 240

c. 250

d. 280

e. None of the above.
Answer #1

2) When the real wage is above the level that equilibrates supply and demand:

a. It creates a deadweight loss in the labor market.

because when real wages is above equilibrium level of demand and supply of labor than the quantity supplied of labor will exceed the quantity demanded of labor which will create dead weight loss in the economy.

3) If Central Bank A cares only about keeping the price level stable and Central Bank B cares only about keeping output at its natural level, then in response to an exogenous increase in the price of oil:

c. Central Bank A should decrease the quantity of money, whereas Central Bank B should increase it.

4) if marginal propensity to save goes up from 0.2 to 0.3 than the total saving will increase by :

B) goes up by 100

5) the level of G needed to make equilibrium Y equal 1,000 is :

e. none of the above

Y = C + I + G + X

C = 100 + 0.6(Y-T)

I = 200

T= 300

Y = 1000

Y = 100 + 0.6 (y-T) + 200 + G

1000 = 100 + 300 + 200 + G

G = 1000 – 600

G = 400